Grade 12th passAlgebra

If one of the roots of the equation x2+(2a+1)x+(a2+2)=0 is twice the value of other than the value of a is??

Shubham
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
suppose rootes are c and d
now
d= 2c
2c + c = – 2a – 1
3c = – 2a – 1
c = (-2a – 1)/3
also
2c^2 = a^2 + 2
(a^2 + 2) / 2 = (-2a- 1/3)^2
solve it
you will get
roots as
c= –3 and d = – 6